OpenStudy (anonymous):

PLEASE HELP a shipping clerk can finish an order in 13 hours, another in 15 hours. If they work together they can finish in how many hours?

OpenStudy (mertsj):

The first clerk does 1/13 of the job in 1 hour. In x hours he does x/13 of the job. the second clerk does 1/15 of the job in 1 hour. In x hours he does x/15 of the job. Working together they do 1 entire job. So: x/13+x/15 = 1

OpenStudy (mertsj):

Solve that and you will get 195/28 hours = 6 27/28 hours

OpenStudy (anonymous):

13 hours - entire work. Work done in ne hour by him = 1/13 And work done in one hour don by other guy = 1/15 Total workdne in oe hour by both of them together = 1/13 + 1/15 = 28/195 Therefore time taken to complete entire work = 195/28 hours.

OpenStudy (anonymous):

thank you guys!

OpenStudy (mertsj):

yw